OpenStudy (anonymous):

Fayed tipped the pizza delivery driver two dollars for pizzas that cost twenty-four dollars. Approximately what percent did he tip the driver? Round your answer to the nearest percent.

jimthompson5910 (jim_thompson5910):

similar example say he tipped $12, that would mean the percentage would be (12/24)*100 = (1/2)*100 = 0.5*100 = 50% This isn't the answer since it's just an example

OpenStudy (anonymous):

12 divided by 4

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

don't get it

jimthompson5910 (jim_thompson5910):

how much did he tip

OpenStudy (anonymous):

$2

jimthompson5910 (jim_thompson5910):

how much is the total bill (ignoring tip)

OpenStudy (anonymous):

22 dollars

jimthompson5910 (jim_thompson5910):

divide the two numbers to get _____

OpenStudy (anonymous):

12

jimthompson5910 (jim_thompson5910):

2/24 is equal to 12?

OpenStudy (anonymous):

oh 0.08333333333

jimthompson5910 (jim_thompson5910):

now multiply that by 100 to convert over to a percentage

OpenStudy (anonymous):

8.33333333

jimthompson5910 (jim_thompson5910):

so it's roughly 8.333% or as a mixed number it is \[\large 8 \frac{1}{3}\%\]

OpenStudy (anonymous):

oh ok thank you

jimthompson5910 (jim_thompson5910):

yw